the possible distance between two tetrahedral voidsin ccp unit cell

Grade:12th pass

2 Answers

Ravleen Kaur
5 years ago
The distance between any two successive (adjacent) tetrahedral void. The angle between body diagonal and an edge is cos inverse of √3a/4, so the projection of the line on an edge is a/4, Similarly the other tetrahedral void will also bea/4 away from its vertex on this edge. So the distance between these two isa−a/4-a/4=a/2
Ravleen Kaur
5 years ago
The distance between any two successive (adjacent) tetrahedral void. The angle between body diagonal and an edge is cos inverse of √3a/4, so the projection of the line on an edge is a/4, Similarly the other tetrahedral void will also bea/4 away from its vertex on this edge. So the distance between these two isa−a/4-a/4=a/2